OpenStudy (anonymous):

A directed line segment from A(-8, -8) to B is divided by P(2, 0) in a ratio 2:1. Where is B? a.) (22,16) b.) (12,8) c.) (7,4)

OpenStudy (anonymous):

Alright so the ratio 2:1 information means the distance from A to P is 2x the distance from P to B, right?

OpenStudy (anonymous):

So there is more than one way to do this

OpenStudy (anonymous):

Okie dokie!

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (anonymous):

Yupp!

OpenStudy (anonymous):

@caseyrt How do I apply the distance formula to this problem? By using the distance formula, I got 2 from those given points.

OpenStudy (anonymous):

ok sweet so first thing to do is find the distance from A to P.

OpenStudy (anonymous):

Okie dokie is 2 correct? =)

OpenStudy (anonymous):

P(2, 0) A(-8, -8) is that right?

OpenStudy (anonymous):

Yes that's right : )

OpenStudy (anonymous):

so our change in x is 10 and change in y is 8. So we should have \[\sqrt{10^2+8^2}\] right?

OpenStudy (anonymous):

yes! \[\sqrt{100+64}\] right?

OpenStudy (anonymous):

yah there we go

OpenStudy (anonymous):

then.... = \[\sqrt{164}\]

OpenStudy (anonymous):

ok so the distance from P to B is going to be 1/2 that

OpenStudy (anonymous):

or \[\sqrt{41}\] because you square the 1/2 when you take it inside the radical

OpenStudy (anonymous):

so set up the distance formula equation from P to B, just use x and y for the B points, and set it equal to sqrt(41)

OpenStudy (anonymous):

All right, makes sense!

OpenStudy (anonymous):

and then you can make an equation for the line you're on in y=mx+b form and sub in the "mx+b" part for y in your expression and solve for x. Then, put x back into the equation you have for your line and solve for y make sense?

OpenStudy (anonymous):

Yes! Wow thank you!

OpenStudy (anonymous):

you bet!